# Concept-as-Tree: Synthetic Data is All You Need for VLM Personalization

### v2

@@ 4.2 Personalized Capability from Synthetic Data
The ability to recognize concepts is fundamental to personalized VLM. As illustrated in Tab. 1, it is evident that after training with extra synthesized data, the model’s recognition ability shows an average improvement of 4.2% on the MC-LLaVA dataset, 2.9% on the Yo’LLaVA dataset, and 2.1% on MyVLM dataset. However, training solely on synthesized data does not yield consistent improvements, which may be attributed to a distribution shift between synthesized data and the real data used for testing. Within the MC-LLaVA framework, despite the use of purely synthesized data, performance still surpasses the baseline, which is attributed to the efficient utilization of visual information relevant to concepts in the MC-LLaVA methods. Including original real images of concepts in the training data can mitigate this phenomenon, resulting in a modest improvement.
We further evaluate the capability of our method to address questions related to personalized concepts. The introduction of synthesized data yields an increase of up to 4.6% in choice-based tasks, 4.7% in Visual Question Answering (VQA) tasks, and 6.2% in captioning tasks, achieving competitive performance with GPT-4o in certain scenarios. Training with purely synthesized data aligns with the trends observed in recognition tasks. Notably, on text-related tasks such as Choice-T and VQA, it consistently enhances model performance. This improvement stems from our targeted modifications in negative sample generation, where specific features of positive samples are altered. By reducing the model’s reliance on these adjusted features, it learns to focus more on the concept’s intrinsic attributes. As a result, the model achieves higher accuracy in concept-related multiple-choice tasks and generates more precise descriptions of the concept’s defining characteristics.
In contrast, RAP-MLLM processes only a single image for each concept. As a result, the synthesized data fails to contribute effectively to its learning, further limiting its comprehension of the concept. While RAP-MLLM may initially outperform models like MyVLM and Yo’LLaVA in vision-based multiple-choice tasks, the synthesized data enables these models to surpass RAP-MLLM by developing a deeper understanding of concept-specific features.
@@ 4.3 Ablations and Analysis
Table 2: Ablation study on tree editing operations. The type and number of operations can both alter the diversity of synthesized data. Excessive or insufficient diversity negatively impacts model performance. TEO = Tree editing operations.
@@
We investigate the impact of tree editing operations on hard negative samples within the MC-LLaVA dataset. Index A represents the Yo’LLaVA baseline without any tree editing operation. The diversity score is defined for a set of data; although there may be a distribution shift between retrieved and generated data, their diversity can still be compared. In Tab. 2, the comparison between A, B, E, and G clearly demonstrates that different tree editing operations lead to variations in the diversity of generated data, resulting in varying task performance. Both Add and Modify operations can enhance diversity to some extent, with Add yielding a greater increase than Modify. In contrast, Remove decreases the diversity of generated results due to a reduction in combinable attributes. In accordance with Sec. 3.1, high-diversity negative samples lead to a more pronounced improvement in tasks related to conversation. However, excessive diversity can introduce noise, thus degrading the performance of the model. For the results of B, C, D, E, F, G and H, we observe that an increase in the number of editing operations per instance amplifies their impact on diversity and results. Therefore, our CaT framework can provide precise control over this process, ensuring that optimal diversity leads to stable improvements.
Figure 6: A significant difference in PCS score between these datasets. Higher PCS score reflects a greater quantity of CS information, correlating with improved synthetic image quality.
@@
Quality of Synthesized Data. As discussed in Sec. 3.3, a higher PCS score indicates that an image contains more CS information, representing a higher-quality personalized sample. From the results shown in Fig. 6, we observe that the distributions of PCS scores for the original positive samples and the synthetic positive samples are quite similar, demonstrating that our synthetic positive samples can effectively serve the same purpose as the original positive samples. Additionally, we find that the proportion of low PCS scores in the original retrieved negative samples is significantly higher than that in our synthetic negative samples, indicating that the quality of synthetic negative samples surpasses that of retrieved negative samples. Based on these two comparisons, we demonstrate that our CaT framework effectively synthesizes high-quality samples, enabling efficient personalization for vision language models.
Figure 7: A comparison of synthesis methods for hard negative samples. Randomly generated prompts, lacking detailed visual descriptions, tend to produce a large number of similar images. Our CaT framework ensures the diversity of synthetic samples, thereby enhancing the model’s personalization capabilities.
@@
Data Component. We analyze the key components of the generated datasets. The original training data consists of provided positive samples and retrieved negative samples. Then we examine the role of synthetic data in these three components by first adding synthetic positive samples and then progressively replacing the retrieved negative samples with synthetic negative samples. At each step, we advance incrementally while keeping the other components unchanged. As shown in Tab. 3, the original Yo’LLaVA already possesses a certain level of personalization capability. However, as samples are progressively transformed into synthesized data, performance on several tasks continues to improve, demonstrating the effectiveness of our CaT approach. The enhancements brought by incorporating easy and hard negative samples vary across different tasks, which is consistent with our observations in Sec. 3.1.
@@
w/ Syn Positive 0.858 (+0.017) 0.816 (+0.015) 0.716 (+0.013) 0.656 (+0.013) 0.721 (+0.020)
w/ Syn Easy Neg 0.873 (+0.015) 0.824 (+0.008) 0.721 (+0.005) 0.667 (+0.011) 0.727 (+0.006)
w/ Syn Hard Neg 0.885 (+0.012) 0.845 (+0.021) 0.738 (+0.017) 0.680 (+0.013) 0.754 (+0.027)
@@
5 Conclusion
In this work, we present a new roadmap to enhance the personalization capability of Vision-Language Models (VLMs) via leveraging synthetic positive and negative images from advanced generative models. Based on our solid and systematic observation experiment, we propose a unified and controllable data synthesis pipeline, which consists of Concept as Tree framework and a well-designed data selection module. Remarkably, our pipeline requires only 1 to 3 user-provided images to work, synthesizing high-quality samples that enable VLMs to perform comparably to models trained on real images in data-scarce scenarios. This pipeline addresses the critical challenge of insufficient user-provided concept data, offering a controllable, interpretable, and data-efficient solution for VLM personalization tasks, paving the way for VLMs to become more accessible and practical human assistants.
Supplementary Material
6 More Implementation Details
Dataset. We utilize datasets from Yo’LLaVA [44] and MyVLM [2]. Yo’LLaVA consists of 40 categories of objects, buildings and people, with 4 to 10 images available per category for training and validation. In contrast, MyVLM encompasses 29 object categories, each containing 7 to 17 images, with 4 images designated for training and the remainder for validation. Additionally, we incorporate the Single concept portion of MC-LLaVA [4] dataset, which is a challenging dataset meticulously constructed from movies, featuring textual descriptions generated with the assistance of GPT-4. MC-LLaVA includes 50 scenarios that encompass various characters and objects, amounting to a total of 118 concepts.
Model Settings. For all LLMs and VLMs involved in the CaT framework, we utilized GPT-4o [1]. After obtaining the concept tree, we synthesize positive and negative samples using the DreamBooth [49] fine-tuned FLUX.1-dev model [32]. The positive samples required for fine-tuning are derived from 1 to 3 original concept images. Finally, we select high-quality images using the CLIP ViT-L/14 visual encoder [46] with our proposed filtering method based on the PCS score. Regarding the personalization of VLMs, we test four methods: MyVLM [2], Yo’LLaVA [44], MC-LLaVA [4] and RAP-MLLM [23]. All configurations adhered to the original papers, and we use LLaVA-1.5-13B [38] as the VLM backbone for all experiments.
Hyperparatemers. The hyperparameters for fine-tuning the FLUX model can be referenced from DreamBooth [49]. In the filtering module, the patch size used in patch shuffle is set to 14, consistent with the CLIP visual encoder we employ. Additionally, based on the distribution of PCS scores from the four datasets, we set the PCS score thresholds for synthesized positive and hard negative samples to 0.3 and 0.1, respectively. For easy negative samples, which do not require CS information, we merely employed conventional text-to-image similarity [46] filtering to ensure that the images matched the prompts.
7 Visualization of Concept Trees
We provide detailed visualizations of Concept Trees, covering three categories: humans in  Figs. 9, 10, 11, 12, 13, 14, 15 and 16, pets in Figs. 17, 18, 19 and 20, and objects in  Figs. 21, 22, 23 and 24. For each concept, we first obtain its concept tree using the CaT. Subsequently, we modify the original tree through three types of tree editing operations: add, modify, and remove. After each operation, a new concept tree is generated, which is then used to synthesize diverse personalized images.
8 CaT Prompt
We provide all the prompts used in the CaT framework. Table 5 includes the three steps for concept tree generation: obtaining the reference image description, performing batch summarization to generate the initial concept tree, and then refining it through self-refinement to obtain the final, well-developed concept tree. Table 6 presents the initialization prompts for the simple negative sample concept tree. Finally, Table 7 provides the three editing operations for the tree—addition, removal, and modification—as well as the prompts used for synthesizing text for image generation based on the concept tree.
9 The Use of PCS Score Filter
@@
We compare our proposed filtering strategy based on the PCS score with the cosine similarity-based filtering strategy that relies on image features. The results presented in Tab. 4 show that the performance of synthetic data without any filtering is even worse than the original Yo’LLaVA [44] baseline. This is because the images synthesized by the generative model may contain low-quality samples that are blurry, unrealistic, or unnatural, which can interfere with the model’s training process. After applying the cosine similarity-based filtering using image features, some low-quality images or those inconsistent with the style of the reference images are filtered out, leading to improved model performance. This demonstrates the necessity of a filtering strategy for synthetic data. However, relying solely on conventional image similarity filtering may fail to remove images that achieve high similarity scores based solely on background, style, or other CA information. Our proposed PCS score-based filtering strategy addresses this limitation by perturbing the synthetic images and calculating the difference in similarity scores before and after perturbation. This approach effectively eliminates the interference of CA information, ultimately selecting images with a higher proportion of CS information. This enables further improvement in the model’s personalization performance in all tasks.
@@
Table 4: Ablation of different filtering methods. The filtering method based solely on cosine similarity cannot exclude images lacking CS information. Our method can ensure high-quality samples, improving model performance.
10 Human Evaluation
@@
The results of the two sets of experiments are presented in Fig. 8. We find that the synthesized positive samples perform comparably to the original positive samples. Meanwhile, the synthesized hard negative samples achieve significantly higher scores than the retrieved hard negative samples. This result clearly demonstrates that our synthesized data can effectively support the personalization training process of vision language models.
Figure 8: Human evaluation of similarity between original and synthesized images compared to reference images: The synthesized positive samples were found to be on par with the original positive samples, while the synthesized hard negative samples demonstrated significantly higher similarity compared to the originally retrieved hard negative samples.
11 Limitation and Future Work
@@
The first limitation stems from our framework’s current inability to handle scenarios involving multiple concepts, primarily due to its dependence on FLUX [32] for image generation. Nevertheless, we believe the inherent scalability of our tree structure paradigm offers promising opportunities for extension through forest-based architectures. Additionally, Future work could explore the integration of enhanced diffusion models with stronger guidance mechanisms, such as layout or box conditions, which have shown potential in better capturing multiple concepts information.
A second limitation arises from potential biases inherited from the LAION-5B [50] dataset used for training FLUX. However, this issue is actively being addressed in ongoing research. Recent studies [51, 18, 54] have made significant progress in developing techniques to enhance fairness and reduce biases in generative models, providing valuable directions for future improvements in our framework.
Finally, the generated concept-centric images may bring data privacy and ethical concerns. This challenge has garnered increasing attention in the research community, with recent works [24, 61, 39] making notable advances in privacy-preserving generation techniques. We anticipate that continued research in GenAI safety will yield effective solutions to mitigate these concerns, enabling more responsible use of synthetic data generation methods.

### v3

@@ 4.2 Personalized Capability from Synthetic Data
The ability to recognize concepts is fundamental to personalized VLM. As illustrated in Tab. 1, it is evident that after training with extra synthesized data, the model’s recognition ability shows an average improvement of 4.1% on the MC-LLaVA dataset, 2.9% on the Yo’LLaVA dataset, and 1.7% on the MyVLM dataset. However, training solely on synthesized data does not yield consistent improvements, which may be attributed to a distribution shift between synthesized data and the real data used for testing. Within the MC-LLaVA framework, despite the use of purely synthesized data, performance still surpasses the baseline, which is attributed to the efficient utilization of visual information relevant to concepts in the MC-LLaVA methods. Including original real images of concepts in the training data can mitigate this phenomenon, resulting in a modest improvement.
We further evaluate the capability of our method to address questions related to personalized concepts. The introduction of synthesized data yields an increase of up to 3.8% in choice-based tasks, 4.2% in Visual Question Answering (VQA) tasks, and 5.3% in captioning tasks, achieving competitive performance with GPT-4o in certain scenarios. Training with purely synthesized data aligns with the trends observed in recognition tasks. Notably, on text-related tasks such as Choice-T and VQA, it consistently enhances model performance. This improvement stems from our targeted modifications in negative sample generation, where specific features of positive samples are altered. By reducing the model’s reliance on these adjusted features, it learns to focus more on the concept’s intrinsic attributes. As a result, the model achieves higher accuracy in concept-related multiple-choice tasks and generates more precise descriptions of the concept’s defining characteristics.
Dataset MC-LLaVA Evaluation (Multi-Concept)
Method/Task Rec Choice-V Choice-T VQA Caption
GPT-4o+Prompt 0.822 0.889 0.680 0.651 0.816
Yo’LLaVA(Real) 0.729 0.602 0.594 0.557 0.611
Yo’LLaVA(Syn) 0.715 0.613 0.579 0.541 0.625
Yo’LLaVA(Real+Syn) 0.732 0.608 0.597 0.569 0.628
Yo’LLaVA(Real+Syn)(P) 0.760 0.634 0.618 0.580 0.647
+3.1% +3.2% +2.4% +2.3% +3.6%
MC-LLaVA(Real) 0.845 0.905 0.695 0.611 0.763
MC-LLaVA(Syn) 0.832 0.918 0.682 0.598 0.776
MC-LLaVA(Real+Syn) 0.849 0.911 0.698 0.624 0.780
MC-LLaVA(Real+Syn)(P) 0.881 0.933 0.714 0.636 0.798
+3.6% +2.8% +1.9% +2.5% +3.5%
RAP-MLLM 0.688 0.690 0.656 0.423 0.748
Table 2: Results of synthetic data used in multi concept scenarios. Our CaT framework also well supports multi-concept scenarios. When trained with a combination of original positive data and all synthetic data, both Yo’LLaVA and MC-LLaVA gain an impressive improvement. The numbers following the “+” symbol represent the improvement of the new SOTA compared to original baseline. Real = Original Data; Syn = Synthetic Data; P = Plus.
Our CaT framework also supports the synthesis of multi-concept personalized training data. By applying attribute constraints across different concept trees within the same scene, we ensure that the synthesized positive samples do not share overlapping attributes. Experimental results in Tab. 2 show that this significantly enhances the capability of multi-concept personalized models.
@@
Table 3: Ablation study on tree editing operations. The type and number of operations can both alter the diversity of synthesized data. TEO = Tree editing operations.
4.3 Ablations and Analysis
@@
Tree Editing Operations. We investigate the impact of tree editing operations on hard negative samples within the MC-LLaVA dataset. Index A represents the Yo’LLaVA baseline without any tree editing operation. The diversity score is defined for a set of data; although there may be a distribution shift between retrieved and generated data, their diversity can still be compared. In Tab. 3, the comparison between A, B, E, and G clearly demonstrates that different tree editing operations lead to variations in the diversity of generated data, resulting in varying task performance. Both Add and Modify operations can enhance diversity to some extent, with Add yielding a greater increase than Modify. In contrast, removing decreases the diversity of generated results due to a reduction in combinable attributes. In accordance with Fig. 3, high-diversity negative samples lead to a more pronounced improvement in tasks related to conversation. However, excessive diversity can introduce noise, thus degrading the performance of the model. For the results of B, C, D, E, F, G and H, we observe that an increase in the number of editing operations per instance amplifies their impact on diversity and results. Therefore, our CaT framework can provide precise control over this process, ensuring that optimal diversity leads to stable improvements.
@@
Quality of Synthesized Data. As discussed in 3.3, a higher PCS score indicates that an image contains more concept-specific information, representing a higher-quality personalized sample. From the results shown in Fig. 6, we observe that the distributions of PCS scores for the original positive samples and the synthetic positive samples are quite similar, demonstrating that our synthetic positive samples can effectively serve the same purpose as the original positive samples. Additionally, we find that the proportion of low PCS scores in the original retrieved negative samples is significantly higher than that in our synthetic negative samples, indicating that the quality of synthetic negative samples surpasses that of retrieved negative samples.
Figure 6: The distribution of PCS scores across different datasets.A higher score indicates that the image contains more concept-specific information.
5 Conclusion
In this work, we present a new roadmap to enhance the personalization capability of Vision-Language Models (VLMs) via leveraging synthetic positive and negative images from advanced generative models. Based on our solid and systematic observation experiment, we propose a unified and controllable data synthesis pipeline, which consists of Concept as Tree framework and a data selection module. Remarkably, our pipeline requires only 1 to 3 user-provided images to work, synthesizing high-quality samples that enable VLMs to perform comparably to models trained on real images in data-scarce scenarios. This pipeline addresses the critical challenge of insufficient user-provided concept data, offering a controllable, interpretable, and data-efficient solution for VLM personalization tasks, paving the way for VLMs to become more practical human assistants.
[…]
Supplementary Material
Appendix A More Implementation Details
Dataset. We utilize datasets from Yo’LLaVA [nguyenyo] and MyVLM [alaluf2025myvlm]. Yo’LLaVA consists of 40 categories of objects, buildings and people, with 4 to 10 images available per category for training and validation. In contrast, MyVLM encompasses 29 object categories, each containing 7 to 17 images, with 4 images designated for training and the remainder for validation. Additionally, we incorporate the Single concept portion of MC-LLaVA [an2024mc] dataset, which is a challenging dataset meticulously constructed from movies, featuring textual descriptions generated with the assistance of GPT-4. MC-LLaVA includes 50 scenarios that encompass various characters and objects, amounting to a total of 118 concepts.
Model Settings. For all LLMs and VLMs involved in the CaT framework, we utilized GPT-4o [achiam2023gpt]. After obtaining the concept tree, we synthesize positive and negative samples using the DreamBooth [ruiz2023dreambooth] fine-tuned FLUX.1-dev model [flux2024]. The positive samples required for fine-tuning are derived from 1 to 3 original concept images. Finally, we select images using the CLIP ViT-L/14 visual encoder [radford2021learning] with our proposed filtering method based on the PCS score. Regarding the personalization of VLMs, we test four methods: MyVLM [alaluf2025myvlm], Yo’LLaVA [nguyenyo], MC-LLaVA [an2024mc] and RAP-MLLM [hao2024rememberretrievegenerateunderstanding]. All configurations adhered to the original papers, and we use LLaVA-1.5-13B [liu2023visual] as the VLM backbone for all experiments.
Hyperparatemers. The hyperparameters for fine-tuning the FLUX model can be referenced from DreamBooth [ruiz2023dreambooth]. In the filtering module, the patch size used in patch shuffle is set to 14, consistent with the CLIP visual encoder we employ. Additionally, based on the distribution of PCS scores from the four datasets, we set the PCS score thresholds for synthesized positive and hard negative samples to 0.3 and 0.1, respectively. For easy negative samples, which do not require CS information, we merely employed conventional text-to-image similarity [radford2021learning] filtering to ensure that the images matched the prompts.
Appendix B More Observation Experiments
The Role of Positive and Negative Samples. We also use MC-LLaVA and MyVLM model to verify the role of positive and negative samples on the MC-LLaVA dataset. The experimental results are in Fig. 7, which are consistent with the conclusions of the Yo’LLaVA model.
Figure 7: More analysis of positive and negative samples.
More Observation Experiment on PCS Score and Diversity. We retrieved and generated negative samples for each concept in Yo’LLaVA dataset. As demonstrated in Tab. 4, our generated hard negative samples demonstrated their high quality in terms of PCS score and diversity measure.
<0.1<0.1 >0.1>0.1 and <0.3<0.3 >0.3>0.3 diversity
Retrieval Negative Samples 49.6% 49.7% 0.7% 0.54
Synthetic Negative Samples 35.9% 63.2% 0.9% 0.68
Table 4: Distribution of PCS score on Yo’LLaVA dataset.
Appendix C More Ablation Experiments
Different Image Encoder. We conducted additional experiments using new visual encoders and VLMs in the filtering process. As shown in Tab. 5, performances between individual encoders are comparable, while combined encoders achieve better results. Furthermore, using stronger VLMs leads to additional performance gains, but brings more costs.
Method / Metric Rec Choice-V Choice-T VQA Caption
CLIP 0.885 0.845 0.738 0.682 0.754
DINOv2 0.897 0.851 0.727 0.679 0.752
Combined (encoders) 0.904 0.854 0.745 0.687 0.760
GPT-4o 0.906 0.862 0.752 0.695 0.765
Table 5: Ablation of different filtering methods, including visual encoders and VLM.
@@ Appendix C More Ablation Experiments
Data Component. We analyze the key components of the generated datasets. The original training data consists of provided positive samples and retrieved negative samples. Then we examine the role of synthetic data in these three components by first adding synthetic positive samples and then progressively replacing the retrieved negative samples with synthetic negative samples. At each step, we advance incrementally while keeping the other components unchanged. As shown in Tab. 6, the original Yo’LLaVA already possesses a certain level of personalization capability. However, as samples are progressively transformed into synthesized data, performance on several tasks continues to improve, demonstrating the effectiveness of our CaT approach. The enhancements brought by incorporating easy and hard negative samples vary across different tasks, which is consistent with our observations in observation experiments.
@@
w/ Syn Positive 0.858 (+1.7%) 0.816 (+1.5%) 0.716 (+1.3%) 0.656 (+1.3%) 0.721 (+2.0%)
w/ Syn Easy Neg 0.873 (+1.5%) 0.824 (+0.8%) 0.721 (+0.5%) 0.667 (+1.1%) 0.727 (+0.6%)
w/ Syn Hard Neg 0.885 (+1.2%) 0.845 (+2.1%) 0.738 (+1.7%) 0.680 (+1.3%) 0.754 (+2.7%)
@@
The Use of PCS Score Metric. We compare our proposed filtering strategy based on the PCS score with the cosine similarity-based filtering strategy that relies on image features. The results presented in Tab. 7 show that the performance of synthetic data without any filtering is even worse than the original Yo’LLaVA [nguyenyo] baseline. This is because the images synthesized by the generative model may contain low-quality samples that are blurry, unrealistic, or unnatural, which can interfere with the model’s training process. After applying the cosine similarity-based filtering using image features, some low-quality images or those inconsistent with the style of the reference images are filtered out, leading to improved model performance. This demonstrates the necessity of a filtering strategy for synthetic data. However, relying solely on conventional image similarity filtering may fail to remove images that achieve high similarity scores based solely on background, style, or other CA information. Our proposed PCS score-based filtering strategy addresses this limitation by perturbing the synthetic images and calculating the difference in similarity scores before and after perturbation. This approach effectively eliminates the interference of CA information, ultimately selecting images with a higher proportion of CS information. This enables further improvement in the model’s personalization performance in all tasks.
@@
Table 7: Ablation of different filtering metrics.
Figure 8: A comparison of hard negative samples.
Appendix D Visualization of Concept Trees
We provide detailed visualizations of Concept Trees, covering three categories: humans in  Figs. 10, 11, 12, 13, 14, 15, 16 and 17, pets in Figs. 18, 19, 20 and 21, and objects in  Figs. 22, 23, 24 and 25. For each concept, we first obtain its concept tree using the CaT. Subsequently, we modify the original tree through three types of tree editing operations: add, modify, and remove. After each operation, a new concept tree is generated, which is then used to synthesize diverse personalized images.
Appendix E CaT Prompt
We provide all the prompts used in the CaT framework. Tab. 8 includes the three steps for concept tree generation: obtaining the reference image description, performing batch summarization to generate the initial concept tree, and then refining it through self-refinement to obtain the final, well-developed concept tree. Tab. 9 presents the initialization prompts for the simple negative sample concept tree. Finally, Tab. 10 provides the three editing operations for the tree—addition, removal, and modification—as well as the prompts used for synthesizing text for image generation.
Appendix F Human Evaluation
@@
The results of the two sets of experiments are presented in Fig. 9. We find that the synthesized positive samples perform comparably to the original positive samples. Meanwhile, the synthesized hard negative samples achieve significantly higher scores than the retrieved hard negative samples. This result clearly demonstrates that our synthesized data can effectively support the personalization training.
Figure 9: Human evaluation of similarity between original and synthesized images compared to reference images: The synthesized positive samples were found to be on par with the original positive samples, while the synthesized hard negative samples demonstrated higher similarity compared to the originally retrieved hard negative samples.
Appendix G Additionally Related Work
Data Quality and Selection. The development of LLMs and VLMs relies heavily on data [dong2025scalable, goyal2024scaling, li2024llm]. The well-known principle is “garbage in, garbage out.” High-quality data can significantly enhance model performance [ridzuan2024review]. To ensure data quality, data selection and filtering are common practices [yi2025bridge, chen2024your]. Primarily, there are rule-based [awotunde2021intrusion] and model-based methods [fabris2025efficient]. Although rule-based methods are simple and heuristic, they are effective. Methods based on LLMs are widely used in data selection. In multimodal scenarios, Clip-filter is one of the most commonly used methods and VLMs are also employed as data filters for self-selection and improvement [wang2024finetuned, hong2024s]. However, since the $\langle\text{sks}\rangle$ we need to evaluate is not in the vocabulary of the language model, retraining the VLM for each new word is cumbersome. Thus, we employ a simple yet effective method by perturbing synthetic images. We evaluate the concept information contained in the original image by calculating the similarity between the image and the reference image before and after perturbation, and apply a predetermined threshold for selection and filtering.
Appendix H Limitation and Future Work
@@
The first limitation stems from our framework’s current inability to handle scenarios involving multiple concepts, primarily due to its dependence on FLUX [flux2024] for image generation. Nevertheless, we believe the inherent scalability of our tree structure paradigm offers promising opportunities for extension through forest-based architectures. Additionally, Future work could explore the integration of enhanced diffusion models with stronger guidance mechanisms, such as layout or box conditions, which have shown potential in capturing multiple concepts information.
A second limitation arises from potential biases inherited from the LAION-5B [schuhmann2022laion5bopenlargescaledataset] dataset used for training FLUX. However, this issue is actively addressed in ongoing research. Recent studies [seshadri2023bias, friedrich2023fair, su2023manifold] have made significant progress in developing techniques to enhance fairness and reduce biases in generative models, providing directions for future improvements in our framework.
Finally, the generated concept-centric images may bring data privacy and ethical concerns. This challenge has garnered increasing attention in the current research community, with recent works [he2024diff, you2024generation, liu2023diffprotect] making notable advances in privacy-preserving generation techniques. We anticipate that continued research in GenAI safety will yield effective solutions to mitigate these concerns, enabling more responsible use of synthetic data generation methods.